DERIVE EXPRESSION FOR EFFICIENCY OF CARNOT CYCLE...

Grade:12

1 Answers

Umakant biswal
7 years ago
@ prathyusha 
 in isothermal expansion , q1= w1 
Q1 = 2.303 log (10 ) v2/v1 RT1 
gas expands at the cost of internal energy 
adiabatic process so, 
w2 = r (t2-t1 ) / y-1 
t1>t2 
in isothermal compression , q3=w3 
q3 = RT22.303 log (10) v3/v4 
at insulating stand w4 = – R (t1-t2)/ y-1 
total wok done = w1+w2+w3+w4 
W = 2.303 log (10) v2/v1 – 2.303 log (10) v3/v4 
W = w1-w3
W= q1-q3 
n = 1- (t2/t1) 
HOPE IT CLEARS YOPUR DOUBT 
ALL THE BEST ..